# D-convolution categories and Hopf algebras

### v1

@@ 3.1.1 Hopf algebras from 1-shifted Lie bialgebras
A 1-shifted metric Lie algebra is a graded Lie algebra $\mathfrak{h}$ with a non-degenerate symmetric invariant bilinear pairing $\kappa$ of degree $1$. From $\kappa$, we may define a new invariant bilinear form $\widetilde{\kappa}(x,y)=\kappa(x,y)(-1)^{|x|}$. This bilinear form is skew-symmetric. Choose a pair of Lagrangian Lie subalgebras $\mathfrak{a}_{\pm}$ of $\mathfrak{g}$, the pairing $\widetilde{\kappa}$ identifies $\mathfrak{a}_{+}$ with the 1-shifted dual of $\mathfrak{a}_{-}$. In particular, this pairing induces cobrackets $\delta_{\pm}$ on $\mathfrak{a}_{\pm}$ which are dual to the brackets of $\mathfrak{a}_{\mp}$ under $\widetilde{\kappa}$. Moreover, $\delta_{\pm}$ are maps of the form
@@
These induces 1-shifted Lie bialgebra structures on $\mathfrak{a}_{\pm}$, as defined in [Pim15] and [NP25]. The corresponding triple $(\mathfrak{h},\mathfrak{a}_{+},\mathfrak{a}_{-})$ is the 1-shifted Manin-triple corresponding to $\mathfrak{a}_{+}$.
In [NP25], it was explained that from such a Manin triple, one can naturally associate a graded Hopf algebra, whose construction we now recall. Let $\{x_{i}\}$ be a set of basis for $\mathfrak{a}_{+}$ with dual basis $\{c^{i}\}$ under $\widetilde{\kappa}$, and let $\mathbf{r}=\sum x_{i}\otimes c^{i}\in\mathfrak{a}_{+}\otimes\mathfrak{a}_{-}$. It turns out that the cobracket $\delta_{\mathfrak{h}}:=(-\delta_{\mathfrak{a}_{+}},\delta_{\mathfrak{a}_{-}})$ is identified with $[\mathbf{r},\Delta]$, where $\Delta$ is the symmetric coproduct of $\mathfrak{h}$. Let $U(\mathfrak{h})$ be the universal enveloping algebra of $\mathfrak{h}$, and let $\mathbb{C}[\delta]$ be a free graded algebra generated by $\delta$ in degree $1$. We denote by $U(\mathfrak{h},\delta):=\mathbb{C}[\delta]\ltimes U(\mathfrak{h})$ the semi-direct product, where
$[\delta,X]=\frac{1}{2}[\nabla\mathbf{r},X],\qquad X\in U(\mathfrak{h}),\nabla:\mathfrak{h}\otimes\mathfrak{h}\to U(\mathfrak{h}).$ (3.2)
@@
Let us now apply this to $\mathfrak{h}=T^{*}[-1]\mathfrak{g}$, the 1-shifted tangent Lie algebra of $\mathfrak{g}$, with Lagrangians and $\mathfrak{a}_{+}=\mathfrak{g},\mathfrak{a}_{-}=\mathfrak{g}^{*}[-1]$. Since $\mathfrak{g}$ is in degree 00, $\mathfrak{h}_{2}=0$ and therefore we have a graded Hopf algebra $U(T^{*}[-1]\mathfrak{g},\delta)/\delta^{2}$. We denote this algebra by $H_{\mathfrak{g}}$. As an algebra, it is generated by $\mathfrak{g},\mathfrak{g}^{*}[-1]$ and $\delta$, whose non-trivial brackets are
$[x_{i},c^{j}]=f_{ki}^{j}c^{k},\qquad[\delta,c^{i}]=-\frac{1}{2}\sum f^{i}_{jk}c^{j}c^{k},\qquad\delta^{2}=0.$ (3.4)
@@ 3.2 Hopf algebras from de-Rham complexes
We need to calculate this more explicitly. The element $\xi_{b}^{ad}$ is the adjoint vector field associated to $b$, which is equal to $-\xi_{b}^{L}+\xi_{b}^{R}$ where $\xi_{b}^{L}$ is the right invariant vector field and $\xi_{b}^{R}$ is the left-invariant vector field. Since $\xi_{b}^{R}$ form a basis for the global sections of vector fields on $G$, we can write
@@
$[\Delta(b),\Delta(c)](g_{1},g_{2})=[1\otimes b,1\otimes c](g_{1},g_{2})+[b\otimes 1,\rho(c)^{op}](g_{1},g_{2}).$ (3.33)
@@
$[\Delta(b),\Delta(c)](g_{1},g_{2})=((1-g_{2}^{-1}g_{1}^{-1})b,c)=\Delta[b,c](g_{1},g_{2}),$ (3.34)
@@
[…]
@@
[…]
@@
[…]
@@
We have defined the algebra $A_{G/G_{ad}}$. One can similarly define algebra $A_{G^{n}/G_{ad}}$, which are simply the semi-direct product of $\mathrm{Dist}(G)\otimes\bigwedge{}^{*}\mathfrak{g}[1]$ with ${\mathcal{A}}_{G}^{\otimes n}$ (except there is only one generator $\delta_{dR}$). One similarly has an equivalence of triangulated categories
@@
[…]
@@
[…]
@@ 4 Quasi-triangular structure and Hopf twist
These results are expected as in Theorem 2.3. However, we note that Theorem 4.1 only implies an abelian equivalence between the centers of the two categories, not the triangulated categories. It is very believable that this equivalence still holds in the triangulated case. We also note that $D_{qs}(D({\mathcal{A}}_{G})\mathrm{-Mod}^{\mathbb{C}^{\times}})$ should be the center of the category ${\mathcal{D}}_{\hbar}(G)^{\mathbb{C}^{\times}}$, which was mis-identified with $D_{qs}(A_{G/G_{ad}}\mathrm{-Mod}^{\mathbb{C}^{\times}})$ in [BZG17].
@@ 4.1 The double of ${\mathcal{A}}_{G}$
Alternatively, one just need to show that $[\Delta(\delta_{dR}),{\mathcal{R}}_{\Omega}]=0$ in $A_{G/G_{ad}}$, which is equivalent to the condition that $(d_{dR}f,\varphi)=-(f,d_{dR}\varphi)$.
@@
[…]
@@ 4.1.1 Hopf subalgebra and character sheaf
The stack $G/G_{ad}$ is the stack of loop spaces $\mathrm{Maps}(S^{1},BG)$ by [BZFN10]. In this work, the authors showed that the DG category $\mathrm{QCoh}(G/G_{ad})$ is the Drinfeld center (in the appropriate DG context) of $\mathrm{QCoh}(BG)=\mathrm{Rep}(G)$. The element ${\mathcal{R}}_{G}$ provides a computable braiding, which can also be expressed in terms of Yetter-Drinfeld data, as in equation (4.16).
@@ 4.2 Double of $H_{G}$
$[b,f]=-\xi_{b}^{L}(f)\delta^{*},\qquad\forall b\in\mathfrak{g}[1],~{}f\in{\mathcal{O}}_{G}.$ (4.24)
@@
This implies that $[b,f]=-\xi_{b}^{L}(f)\delta^{*}$ as desired.
@@
and we find $[\delta,f]=[\delta_{dR},f]$. Similarly, the double coproduct of $b$ is of the form
@@ 4.2.1 Shifted Poisson Lie groups and deformations
$[b,f]=-\xi_{b}^{L}(f)\delta^{*}_{dR},\qquad b\in\mathfrak{g}[1],f\in{\mathcal{O}}_{G}.$ (4.37)
[…]
$[b,f]=-2\xi_{b}^{L}(f)\delta^{*}=-2\{\xi_{b}^{L},f\}\delta^{*},$ (4.38)
@@ 4.3 Identifying the doubles
or equivalently, $[{\mathcal{R}}_{G}^{21,-1},1\otimes\delta_{dR}]{\mathcal{R}}_{G}^{21}=\sum x_{i}\otimes c^{i}$. Let us write ${\mathcal{R}}_{G}={\mathcal{R}}^{(1)}\otimes{\mathcal{R}}^{(2)}$, then
$[{\mathcal{R}}_{G}^{21,-1},1\otimes\delta_{dR}]{\mathcal{R}}_{G}^{21}=-S({\mathcal{R}}^{(2)}){\mathcal{R}}^{(2)^{\prime}}\otimes\delta_{dR}({\mathcal{R}}^{(1)}){\mathcal{R}}^{(1)^{\prime}}=-S({\mathcal{R}}^{(2)}){\mathcal{R}}^{(2)^{\prime}}\otimes x_{i}({\mathcal{R}}^{(1)}){\mathcal{R}}^{(1)^{\prime}}c^{i}.$ (4.43)
The element $-S({\mathcal{R}}^{(2)}){\mathcal{R}}^{(2)^{\prime}}\otimes x_{i}({\mathcal{R}}^{(1)}){\mathcal{R}}^{(1)^{\prime}}$ belongs to $D({\mathcal{A}}_{G})\otimes{\mathcal{O}}_{G}$, we pair them with $g\in G$, which gives
$-S(gx_{i})g=x_{i}g^{-1}g=x_{i}.$ (4.44)
@@
$-S({\mathcal{R}}^{(2)}){\mathcal{R}}^{(2)^{\prime}}\otimes x_{i}({\mathcal{R}}^{(1)}){\mathcal{R}}^{(1)^{\prime}}=\sum x_{i}\otimes c^{i},$ (4.45)

### v2

@@ 3.1.1 Hopf algebras from 1-shifted Lie bialgebras
A 1-shifted metric Lie algebra is a graded Lie algebra $\mathfrak{h}$ with a non-degenerate symmetric invariant bilinear pairing $\kappa$ of degree $1$. From $\kappa$, we may define a new invariant bilinear form $\widetilde{\kappa}(x,y)=\kappa(x,y)(-1)^{|x|}$. This bilinear form is skew-symmetric. Choose a pair of Lagrangian Lie subalgebras $\mathfrak{a}_{\pm}$ of $\mathfrak{h}$, the pairing $\widetilde{\kappa}$ identifies $\mathfrak{a}_{+}$ with the 1-shifted dual of $\mathfrak{a}_{-}$. In particular, this pairing induces cobrackets $\delta_{\pm}$ on $\mathfrak{a}_{\pm}$ which are dual to the brackets of $\mathfrak{a}_{\mp}$ under $\widetilde{\kappa}$. Moreover, $\delta_{\pm}$ are maps of the form
@@
The following was shown in [NP25, Proposition 2.3].
Proposition 3.1.
The cobrackets $\delta_{\pm}$ satisfies
$\delta_{\pm}\otimes 1+1\otimes\delta_{\pm}(\delta_{\pm})=0\in\mathrm{Sym}^{3}(\mathfrak{a}_{\pm}),\qquad\delta_{\pm}([X,Y])=[\delta_{\pm}(X),\Delta(Y)]+(-1)^{|X|}[\Delta(X),\delta_{\pm}(Y)].$ (3.2)
In [PIM15] and [NP25], a Lie algebra $\mathfrak{a}$ together with a cobracket $\delta$ satisfying equation (3.2) is called a 1-shifted Lie bialgebra, and the triple $(\mathfrak{h},\mathfrak{a}_{+},\mathfrak{a}_{-})$ is called the Manin-trimple corresponding to $\mathfrak{a}_{+}$. It was shown in [NP25, Theorem 2.5] that a 1-shifted Lie bialgebra $\mathfrak{a}$ gives rise to a canonical Manin-triple $(D(\mathfrak{a}),\mathfrak{a},\mathfrak{a}^{*}[-1])$, and vice versa.
In [NP25], it was explained that from such a Manin triple, one can naturally associate a graded Hopf algebra, whose construction we now recall. Let $\{x_{i}\}$ be a set of basis for $\mathfrak{a}_{+}$ with dual basis $\{c^{i}\}$ under $\widetilde{\kappa}$, and let $\mathbf{r}=\sum x_{i}\otimes c^{i}\in\mathfrak{a}_{+}\otimes\mathfrak{a}_{-}$. It turns out that the cobracket $\delta_{\mathfrak{h}}:=(-\delta_{\mathfrak{a}_{+}},\delta_{\mathfrak{a}_{-}})$ is identified with $\mathbf{r},\Delta]$, where $\Delta$ is the symmetric coproduct of $\mathfrak{h}$. Let $U(\mathfrak{h})$ be the universal enveloping algebra of $\mathfrak{h}$, and let $\mathbb{C}[\delta]$ be a free graded algebra generated by $\delta$ in degree $1$. We denote by $U(\mathfrak{h},\delta):=\mathbb{C}[\delta]\ltimes U(\mathfrak{h})$ the semi-direct product, where
$\delta,X]=\frac{1}{2}[\nabla\mathbf{r},X],\qquad X\in U(\mathfrak{h}),\nabla:\mathfrak{h}\otimes\mathfrak{h}\to U(\mathfrak{h}).$ (3.3)
@@
Let us now apply this to $\mathfrak{h}=T^{*}[-1]\mathfrak{g}$, the 1-shifted tangent Lie algebra of $\mathfrak{g}$, with Lagrangians and $\mathfrak{a}_{+}=\mathfrak{g},\mathfrak{a}_{-}=\mathfrak{g}^{*}[-1]$. Since $\mathfrak{g}$ is in degree 0, $\mathfrak{h}_{2}=0$ and therefore we have a graded Hopf algebra $U(T^{*}[-1]\mathfrak{g},\delta)/\delta^{2}$. We denote this algebra by $H_{\mathfrak{g}}$. As an algebra, it is generated by $\mathfrak{g},\mathfrak{g}^{*}[-1]$ and $\delta$, whose non-trivial brackets are
$x_{i},c^{j}]=f_{ki}^{j}c^{k},\qquad[\delta,c^{i}]=-\frac{1}{2}\sum f^{i}_{jk}c^{j}c^{k},\qquad\delta^{2}=0.$ (3.5)
@@ 3.2 Hopf algebras from de-Rham complexes
Proof.
First, (3.25) indeed defines a Hopf algebra structure on ${\mathcal{A}}_{G}$. The fact its restriction on $\Omega_{G}$ is a Hopf algebra structure follows from the fact that $\Omega_{G}$ is identified with functions on the (DG) group $T[1]G$, and equation (3.25) is simply the corresponding map on the algebra of functions. The fact that it extends to the generator $\delta_{dR}$ follows from the fact that for any map $h:X\to Y$, one has $h^{*}(d_{dR}\omega)=d_{dR}h^{*}(\omega)$ for any $\omega\in\Omega(X)$; therefore $m_{G}^{*}([\delta_{dR},-])=[\delta_{dR}\otimes 1+1\otimes\delta_{dR},m_{G}^{*}(-)]$.
We are left to show that the equivalence of Proposition 2.4 is one of monoidal categories. To this end, we have the following diagram of categories
$C({\mathcal{A}}_{G}\mathrm{-Mod}^{\mathbb{C}^{\times}})\times C({\mathcal{A}}_{G}\mathrm{-Mod}^{\mathbb{C}^{\times}})$$C({\mathcal{A}}_{G\times G}\mathrm{-Mod}^{\mathbb{C}^{\times}})$$C({\mathcal{A}}_{G}\mathrm{-Mod}^{\mathbb{C}^{\times}})$$C({\mathcal{M}}_{\hbar}(G)^{\mathbb{C}^{\times}})\times C({\mathcal{M}}_{\hbar}(G)^{\mathbb{C}^{\times}})$$C({\mathcal{M}}_{\hbar}(G\times G)^{\mathbb{C}^{\times}})$$C({\mathcal{M}}_{\hbar}(G)^{\mathbb{C}^{\times}})$$\scriptstyle{\varphi}$$\scriptstyle{D_{\hbar}\times D_{\hbar}}$$\scriptstyle{\psi}$$\scriptstyle{D_{\hbar}}$$\scriptstyle{D_{\hbar}}$$\scriptstyle{\hbar_{1}=\hbar_{2}}$$\scriptstyle{m_{G,*}}$ (3.27)
The horizontal diagrams define the monoidal product on the categories involved. The functor $\varphi$ is pullback from ${\mathcal{A}}_{G}\otimes{\mathcal{A}}_{G}$ to ${\mathcal{A}}_{G\times G}$ which is identity on $\Omega_{G\times G}$ and maps $\delta_{dR,G\times G}$ to $\delta_{dR}\otimes 1+1\otimes\delta_{dR}$, and the functor $\psi$ is the pullback functor along the coproduct of $\Omega_{G}$. On the second diagram, one first take the tensor product of modules over $\mathbb{C}[\![\hbar]\!]$, then pushforward along $m_{G}$. By Proposition 2.4, the second square commutes, we just need to verify that the first diagram commutes.
Recall from equation (2.10), the functor $D_{\hbar}$ sends $N\in C({\mathcal{A}}_{G}\mathrm{-Mod}^{\mathbb{C}^{\times}})$ to $D_{G}^{\hbar}\otimes_{{\mathcal{O}}_{X}}N$ with Koszul differential. Therefore, we have
$D_{\hbar}(N_{1})\otimes_{\mathbb{C}[\![\hbar]\!]}D_{\hbar}(N_{2})=(D_{G}^{\hbar}\otimes_{\mathbb{C}[\![\hbar]\!]}D_{G}^{\hbar})\otimes_{{\mathcal{O}}_{G\times G}}(N_{1}\otimes N_{2})=D_{G\times G}^{\hbar}\otimes_{{\mathcal{O}}_{G\times G}}(N_{1}\otimes N_{2}),$ (3.28)
with differential
$\sum v_{i}^{(1)}\otimes\omega^{i,(1)}+v_{i}^{(2)}\otimes\omega^{i,(2)}+\hbar\otimes(\delta_{dR}^{(1)}+\delta_{dR}^{(2)})+d_{N_{1}}+d_{N_{2}},$ (3.29)
where the supscripts $(1),(2)$ refer to the action on the first and second factor of $D_{G}^{\hbar}$ (and $N_{1}\otimes N_{2}$). The above DG module coincides with $D_{\hbar}\varphi(N_{1}\otimes N_{2})$.
The commutativity of the diagram gives the monoidal isomorphism for the functor $D_{\hbar}$, and the fact that it is associative is clear. The descent of this monoidal functor to $D_{qs}$ gives the desired equivalence at the level of monoidal categories.
@@
We need to calculate this more explicitly. The element $\xi_{b}^{ad}$ is the adjoint vector field associated to $b$, which is equal to $\xi_{b}^{L}+\xi_{b}^{R}$ where $\xi_{b}^{L}$ is the right invariant vector field and $\xi_{b}^{R}$ is the left-invariant vector field. Since $\xi_{b}^{R}$ form a basis for the global sections of vector fields on $G$, we can write
@@
$\Delta(b),\Delta(c)](g_{1},g_{2})=[1\otimes b,1\otimes c](g_{1},g_{2})+[b\otimes 1,\rho(c)^{op}](g_{1},g_{2}).$ (3.37)
@@
$\Delta(b),\Delta(c)](g_{1},g_{2})=((1-g_{2}^{-1}g_{1}^{-1})b,c)=\Delta[b,c](g_{1},g_{2}),$ (3.38)
@@
$(G_{L}/G_{\Delta})^{2}$$G_{L}^{2}/G_{\Delta}$$G_{L}/G_{\Delta}$$(G_{\Delta}\setminus G^{2}/G_{\Delta})^{2}$$G_{\Delta}\setminus G^{2}\times_{G_{\Delta}}G^{2}/G_{\Delta}$$G_{\Delta}\setminus G^{2}/G_{\Delta}$$\scriptstyle{p}$$\scriptstyle{m}$$\scriptstyle{\widetilde{m}}$$\scriptstyle{\widetilde{p}}$ (3.41)
@@
$G_{\Delta}\setminus G^{2}\times_{G_{\Delta}}G^{2}\times_{G_{\Delta}}G^{2}/G_{\Delta}$$G_{\Delta}\setminus G^{2}\times_{G_{\Delta}}G^{2}/G_{\Delta}$$G_{\Delta}\setminus G^{2}\times_{G_{\Delta}}G^{2}/G_{\Delta}\times G_{\Delta}\setminus G^{2}/G_{\Delta}$$(G_{\Delta}\setminus G^{2}/G_{\Delta})^{2}$$\scriptstyle{\widetilde{m}_{12}}$$\scriptstyle{\widetilde{p}}$$\scriptstyle{p}$$\scriptstyle{m_{12}}$ (3.45)
@@
$G_{L}^{3}/G_{\Delta}$$G_{L}^{2}/G_{\Delta}$$G_{L}^{2}/G_{\Delta}\times G_{L}/G_{\Delta}$$(G_{L}/G_{\Delta})^{2}$$\scriptstyle{\widetilde{m}_{12}}$$\scriptstyle{\widetilde{p}}$$\scriptstyle{p}$$\scriptstyle{m_{12}}$ (3.46)
@@
We have defined the algebra $A_{G/G_{ad}}$. One can similarly define algebra $A_{G^{n}/G_{ad}}$, which are simply the semi-direct product of $\mathrm{Dist}(G)\otimes\bigwedge{}^{*}\mathfrak{g}[1]$ with ${\mathcal{A}}_{G}^{\otimes n}$ (except there is only one generator $\delta_{dR}$, whose coproduct can be similarly treated as in Corollary 3.9). One similarly has an equivalence of triangulated categories
@@
$G_{L}^{3}/G_{\Delta}$$G_{L}^{2}/G_{\Delta}$$G_{L}^{2}/G_{\Delta}\times G_{L}/G_{\Delta}$$(G_{L}/G_{\Delta})^{2}$$\scriptstyle{\widetilde{m}_{12}}$$\scriptstyle{\widetilde{p}}$$\scriptstyle{p}$$\scriptstyle{m_{12}}$ (3.50)
@@
$G_{\Delta}\times G_{L}^{3}$$G_{\Delta}\times G_{L}^{2}$$G_{L}^{2}\times G_{L}$$(G_{L})^{2}$$\scriptstyle{\widetilde{m}_{12}}$$\scriptstyle{\widetilde{p}}$$\scriptstyle{p}$$\scriptstyle{m_{12}}$ (3.51)
@@ 4 Quasi-triangular structure and Hopf twist
These results are expected as in Theorem 2.3. However, we note that Theorem 4.1 only implies an abelian equivalence between the centers of the two categories, not the triangulated categories. It is very believable that this equivalence still holds in the triangulated case. We also note that $D_{qs}(D({\mathcal{A}}_{G})\mathrm{-Mod}^{\mathbb{C}^{\times}})$ should be the center of the category ${\mathcal{D}}_{\hbar}(G)^{\mathbb{C}^{\times}}$, which was mis-identified with $D_{qs}(A_{G/G_{ad}}\mathrm{-Mod}^{\mathbb{C}^{\times}})$ in [BG17].
@@ 4.1 The double of ${\mathcal{A}}_{G}$
Alternatively, one just need to show that $\Delta(\delta_{dR}),{\mathcal{R}}_{\Omega}]=0$ in $A_{G/G_{ad}}$, which is equivalent to the condition that $(d_{dR}f,\varphi)=-(f,d_{dR}\varphi)$.
@@
$M{{\mathcal{O}}_{G}\otimes M}$${\mathcal{O}}_{G}\otimes M$$M\scriptstyle{\rho_{M}}$$\scriptstyle{S\otimes 1}$$\scriptstyle{\mathrm{act}_{M}}$ (4.17)
@@ 4.1.1 Hopf subalgebra and character sheaf
The stack $G/G_{ad}$ is the stack of loop spaces $\mathrm{Maps}(S^{1},BG)$ by [BFN10]. In this work, the authors showed that the DG category $\mathrm{QCoh}(G/G_{ad})$ is the Drinfeld center (in the appropriate DG context) of $\mathrm{QCoh}(BG)=\mathrm{Rep}(G)$. The element ${\mathcal{R}}_{G}$ provides a computable braiding, which can also be expressed in terms of Yetter-Drinfeld data, as in equation (4.16).
@@ 4.2 Double of $H_{G}$
$b,f]=-\xi_{b}^{L}(f)\delta^{*},\qquad\forall b\in\mathfrak{g}[1],~f\in{\mathcal{O}}_{G}.$ (4.24)
@@
This implies that $b,f]=-\xi_{b}^{L}(f)\delta^{*}$ as desired.
@@
and we find $\delta,f]=[\delta_{dR},f]$. Similarly, the double coproduct of $b$ is of the form
@@ 4.2.1 Shifted Poisson Lie groups and deformations
$b,f]=-\xi_{b}^{L}(f)\delta^{*}_{dR},\qquad b\in\mathfrak{g}[1],f\in{\mathcal{O}}_{G}.$ (4.37)
Note that if we set $\delta^{*}_{dR}=0$, then the quotient $H^{*}_{G}/(\delta^{*}_{dR})$ can be identified with functions on $T^{*}[-1]G$. This is in fact an identification of Hopf algebras, if we identify $T^{*}[-1]G=G\ltimes\mathfrak{g}^{*}[-1]$ as the group extension of $G$ by $\mathfrak{g}^{*}[-1]$. It’s Lie algebra is precisely $T^{*}[-1]\mathfrak{g}$, which was used to construct $H_{\mathfrak{g}}$. The 1-shifted Lie bialgebra structure on $T^{*}[-1]\mathfrak{g}$ gives rise to a 1-shfited Poisson structure $\{-,-\}$ on $T^{*}[-1]G$, making it a 1-shifted Poisson Lie group. This is very similar to the fact that a Lie bialgebra structure on $\mathfrak{g}$ is related to a Poisson Lie group structure on $G$. We see that $H^{*}_{G}$ can be viewed as a quantization, in the sense that
$b,f]=-2\xi_{b}^{L}(f)\delta^{*}=-2\{\xi_{b}^{L},f\}\delta^{*},$ (4.38)
@@ 4.3 Identifying the doubles
or equivalently, ${\mathcal{R}}_{G}^{21,-1},1\otimes\delta_{dR}]{\mathcal{R}}_{G}^{21}=\sum x_{i}\otimes c^{i}$. Let us write ${\mathcal{R}}_{G}={\mathcal{R}}^{(1)}\otimes{\mathcal{R}}^{(2)}$, then
${\mathcal{R}}_{G}^{21,-1},1\otimes\delta_{dR}]{\mathcal{R}}_{G}^{21}=-S({\mathcal{R}}^{(2)}){\mathcal{R}}^{(2)^{\prime}}\otimes\delta_{dR}({\mathcal{R}}^{(1)}){\mathcal{R}}^{(1)^{\prime}}=-S({\mathcal{R}}^{(2)}){\mathcal{R}}^{(2)^{\prime}}\otimes x_{i}({\mathcal{R}}^{(1)}){\mathcal{R}}^{(1)^{\prime}}c^{i}.$ (4.43)
The element $S({\mathcal{R}}^{(2)}){\mathcal{R}}^{(2)^{\prime}}\otimes x_{i}({\mathcal{R}}^{(1)}){\mathcal{R}}^{(1)^{\prime}}$ belongs to $D({\mathcal{A}}_{G})\otimes{\mathcal{O}}_{G}$, we pair them with $g\in G$, which gives
$S(gx_{i})g=x_{i}g^{-1}g=x_{i}.$ (4.44)
@@
$S({\mathcal{R}}^{(2)}){\mathcal{R}}^{(2)^{\prime}}\otimes x_{i}({\mathcal{R}}^{(1)}){\mathcal{R}}^{(1)^{\prime}}=\sum x_{i}\otimes c^{i},$ (4.45)